# BlockLLM: Multi-tenant Finer-grained Serving for Large Language Models

### v1

@@ Abstract
The growing demand for Large Language Models (LLMs) across diverse applications has prompted a paradigm shift in the design of deep learning serving systems. Deploying LLMs, especially in multi-tenant environments, presents considerable challenges due to their high computational and memory demands. We present BlockLLM, a serving system that exploits the potential of sharing components among fine-tuned LLM models to offer an efficient and flexible solution for LLM workloads. BlockLLM partitions the models into finer-grained blocks to enable the reuse of model components and independent provisioning to improve the computation efficiency. BlockLLM consists of an offline block zoo, for storing the blocks, and an online system to serve the requests through chains of blocks. It offers multi-fold flexibility: (1) Adaptive assembly of block chains on-the-fly is achieved with the help of equivalence evaluation among blocks in the zoo. (2) We enable per-block batch size and configure best-effort KV cache coordination at individual block level. (3) We adopt speculative execution and locality-aware block placement to mitigate the communication costs from dynamic block resource allocation. Our evaluation demonstrates that BlockLLM reduces memory and storage footprints and improves computation efficiency, outperforming existing serving approach in 95%ile latency and GPU utilization by 33.5% and 20.1%, respectively.
@@ 1 Introduction
The rise of Large Language Models (LLMs) marks a transformative milestone in deep learning. Their unprecedented abilities in natural language processing [28], from language translation [6] to question-answering [1] and complex reasoning [20, 40], are reshaping technological frontiers. Yet, the strength of LLMs—rooted in their vast parameter spaces—comes at the price of significant computation costs [5]. The deployment of LLMs is an arduous endeavor, often necessitating the use of hundreds or even thousands of extremely expensive computation devices like GPUs, thus imposing great challenges on unleashing LLM’s full potential for more people [36, 3].
@@
With the growing reliance on LLMs, fine-tuning has emerged as a critical technique to efficiently adapt foundation models to handle various tasks across different domains [16, 9]. Fine-tuning updates existing parameters or introduces additional parameters using domain-specific data. As LLMs become ubiquitous, serving these fine-tuned models from a multi-tenant cluster, whether in a private cloud for first-party services or a public cloud for third-party users, is starting to emerge as an important consideration.
@@
Recent serving systems and approaches focus on model-internal optimization, such as the acceleration of matrix computations [8, 2] and minimizing device memory fragmentation [18] to enable parallel decoding to circumvent the limitations imposed by auto-regressive models that process a single token at a time [11, 26]. While these advancements significantly accelerate the inference process for individual LLMs, the challenge remains to serve multiple tenants’ fine-tuned models in a shared cluster. Each LLM requires substantial resources, and the hardware cost makes it prohibitive to liberally expand the cluster to add dedicated resources for each tenant. Thus, the question is: how can LLM serving meet the latency performance goals of tenants’ fine-tuned models while also ensuring high throughput and optimal utilization cluster-wide?
@@
We introduce BlockLLM, an efficient and flexible serving system tailored for LLM workloads in a multi-tenant cluster. BlockLLM enables LLM model component reuse by breaking down an LLM into smaller and shareable components which we call “blocks”. The central observation forming the basis for BlockLLM is that LLM fine-tuning offers the opportunity for sharing. Parameter-efficient fine-tuned models have an inherent modular architecture where parameters from specific layers are added or altered. Our observations with full-parameter fine-tuned models indicate that certain model components produce outputs with high similarity, suggesting their potential for sharing as well. Block sharing reduces both memory and storage demands. Furthermore, by reallocating memory - previously consumed by redundant parameters - toward supporting larger data batch size, BlockLLM achieves higher computation efficiency and overall throughput.
@@
BlockLLM consists of an offline “block zoo”—a repository of LLM blocks, and an online serving system capable of dynamically scheduling blocks to serve various applications (Figure 1). BlockLLM’s practice brings three opportunities. (1) BlockLLM can support adaptive serving, which enables on-the-fly assembly of blocks for each request, moving beyond the traditional static chain of blocks. (2) Each block can be individually configured, allowing for customized batch sizes and autonomous handling of queued requests. (3) BlockLLM enables dynamic resource allocation, with each block capable of independently scaling and being placed without the constraints of model-level boundaries.
@@
First, to ascertain the equivalence of blocks within BlockLLM, we consider two methods subject to block architecture. For blocks of the same architecture that naturally support seamless request routing, cosine similarity metrics are utilized to measure parametric differences. For blocks of disparate embedding sizes, the challenge is more intricate. We compare the output distribution of vocabulary probabilities,
which serves as an indicator for functional equivalence between transformer layers. Inspired by the techniques of merging computer vision models [30], we design a generalizable stitching block as an intermediary to route requests among these blocks so that the chain of blocks to serve each request can be adaptively adjusted online.
Second, to manage the KV cache generated by different requests, BlockLLM implements a principled approach to dispatch requests. Per-block configuration introduces complexity to the handling of the KV cache, particularly as requests might be allocated to devices that do not currently hold the corresponding cache, necessitating either cache communication or recalculation. We analyze the trade-off between the recalculation and new IO costs and use it to compute different priorities for different candidate block instances, including the blocks specified in the chain and its equivalent blocks. We adopt a best-effort KV coordination policy where devices possessing the KV cache are prioritized when dispatching the request.
Third, to address the inherent latency costs posed by inter-block communication, BlockLLM integrates speculative execution for identified bottleneck blocks, thereby processing potential request pathways ahead-of-time. We use surrogate models to predict the outputs of blocks, facilitating computation ahead of actual output. In parallel, we implement a locality-aware strategy for block placement, putting interdependent blocks nearby, preferably within the same server. This approach reduces the frequency of costly inter-server communications and leverages the high-speed intra-server connections to expedite data transfer and reduce inference latency.
@@
  • We leverage the LLM fine-tuning characteristic and demonstrate the effectiveness of finer-grained LLM serving in terms of reducing memory and storage requirements and optimizing computation resource utilization.
  • We construct a block zoo that stores LLMs in finer granularity to enable block reuse and establishes the relationship among these blocks to facilitate adaptive online serving.
@@
  • We implement and evaluate BlockLLM, a multi-tenant serving system. We evaluate BlockLLM in a 12-A100 cluster. BlockLLM achieves comparable median latency with traditional per-model provisioning and reduces the 95%ile latency by 33.5%. The overall average throughput is elevated by 1.71x. GPU SM efficiency is improved by 20.1%.
@@ 2.1 Background
LLM serving. LLM-based applications have grown to be one of the major workloads in GPU clusters. Given their colossal scale, the burden of serving these models has emerged as a critical issue. In the offline stage, large storage is required to store the model checkpoints and engines. In the online stage, numerous powerful computation units (e.g., GPU devices) with large memory spaces are necessary to load and execute the models [35, 31]. LLMs’ unique auto-regressive nature necessitates caching KV matrices of tokens on devices to avoid redundant recalculations [32]. Every request has its own KV cache, which progressively grows in length as the auto-regressive process continues. Moreover, when models are inevitably deployed in a distributed manner, a high-capacity network is indispensable to transfer the intermediate activations. Such a burden is particularly pronounced in multi-tenant clusters, which cater to diverse LLM applications, each with its dedicated models and performance objectives.
@@
2.2 Finer-grained Serving
Key idea. To enhance the efficiency of LLM serving systems, we advocate a finer-grained approach. Rather than treating each model as an indivisible unit, we propose BlockLLM, which partitions LLMs into smaller components provisioned independently. These components are envisioned as “blocks”. We could maintain only one copy of a block even if it is used by multiple models. The immediate benefit is the decreased demand for memory and storage.
@@ 2.2 Finer-grained Serving
PEFT models are particularly well-suited to this approach due to their inherent modular architecture. For example, as illustrated in Figure 2, two models could be broken down into five blocks, with blue blocks being common components shared between them. Moreover, our analysis of FPFT models indicates that blocks with varied parameters often produce highly similar outputs. Figure 3(a) presents the parameter cosine similarity of each Transformer layer between LLaMA and Vicuna/Alpaca (FF from LLaMA). The average similarity is 0.9927. This observation underpins the feasibility of reusing blocks across different FPFT models, even when parameter values differ. Enabling the reuse of model components would improve resource efficiency by letting device memory hold more critical data.
(a) LLaMA 7B and Vicuna 7B.
(b) LLaMA 7B and Alpaca 7B.
Figure 3: Cosine similarity of parameters in each Transformer layer of two models.
@@
O3: Dynamic resource allocation. BlockLLM enhances resource allocation with its ability to independently scale each block. Blocks that are frequently accessed or computationally intensive can dynamically be allocated with more resources. It also liberates block placement from the constraints associated with monolithic model architectures. However, it does introduce a challenge in the form of increased communication costs. Effectively mitigating these costs is crucial to achieving truly versatile resource allocation within BlockLLM (5.2, 5.3).
@@ 2.3 Comparison with Existing Work
Existing serving systems. Existing serving systems for LLM workload follow the practice of serving traditional DNNs. Each model operates within a dedicated engine and is provisioned as a monolithic unit. An alternative approach is multi-tasking. A typical example is S-LoRA [34], which exploits memory paging to pack multiple LoRAs within one engine. Another line of optimizations focuses on the model internals—like accelerating Attention layer computations (FlashAttention [8]), managing KV cache (PagedAttention [18] in vLLM), and efficiently scheduling varying sequence lengths of requests (Orca [42]). These efforts are complementary to BlockLLM. BlockLLM tackles the inefficiency of serving LLM workloads in a multi-tenant cluster by leveraging the characteristics of LLM fine-tuning. Block-granularity provisioning improves the serving efficiency from two perspectives.
@@
Table 1: Percentage of shared parameters of different PEFT techniques.
@@ 3 BlockLLM Design
We provide an overview of BlockLLM and explain its three design challenges.
@@ 3.1 System Overview
Workflow. During serving, the workflow is as follows. When a new request arrives, the scheduler assigns a chain of blocks based on the application it belongs to. The scheduler initially determines whether there is an available instance of the first block in the chain. If available, the request is forwarded to the target block instance. Otherwise, the scheduler either identifies a suitable alternative among the existing equivalent block instances or deploys a new block instance on an available device. Upon receipt of the request, the device’s agent directs it to the designated block instance for processing. Then, once the execution is completed, the agent is responsible for directing the request through the subsequent block in the chain. The process ends with the completion of the request signified by an EOS token, whereupon the final agent relays the output back to the scheduler, thus concluding the response cycle of the request.
@@ 3.2 Design Challenges
We explain three design challenges in BlockLLM, which are briefly mentioned in 2.2.
@@
C3: Mitigate communication overhead. Versatile resource allocation of finer-grained blocks provides several benefits for online serving, as mentioned in 2.2. Yet BlockLLM incurs additional non-negligible communication overhead. Specifically, each time a request batch is directed to a different device, it necessitates intra-server or inter-server communication. This communication might act as a bottleneck for the batch, directly contributing to increased latency. The communication overhead also linearly grows with the number of block instances passed by a request batch during its lifecycle. Furthermore, the volume of data transferred escalates with the use of larger batch sizes per block. Figure 7 breaks down the latency components involved in generating one new token across various request batches. Without any optimization to the communication, for a request batch of size 32 and length 512, the actual computation constitutes 62.9% of the total latency, whereas the inter-server communication accounts for 36.4%.
@@ 4 Block Zoo
  • How to handle request routing between two equivalent blocks with different embedding sizes? (C1, 4.3)
@@ 4.1 Equivalence
PE models inherently possess equivalent parameters for reusing. We thus focus on verifying if equivalence exists among FF model components. Our goal is to find such equivalence so that some of the FF model components can be shared as well once they are partitioned into blocks. We consider two scenarios here.
@@
Models with an identical architecture. First, when two models share an identical architecture, we can establish the equivalence to optimize request routing. For instance, if tail layers (Layer $i$ to $L$) are determined to be equivalent, as depicted in Figure 9, we have the flexibility to route requests from head layers of Model B to tail layers of Model A, which becomes particularly useful when the tail layers of Model B is experiencing high load. This strategy is also beneficial in conserving resources and reducing loading overhead when no instance of the tail layer of Model B is actively deployed.
In this case, the degree of equivalence between model components can be systematically quantified. In BlockLLM, we employ the canonical metric, cosine similarity, to evaluate the parametric similarity of each Transformer layer. The similarity is defined as the weighted average of cosine similarities of all constituent parameters within each Transformer layer. That is, for layer $A_{i}$ and $B_{i}$ is:
$\vspace{-1mm}Eq(A_{i},B_{i})=\sum_{p=1}^{n}(s(A_{i}^{p})\times cos(A_{i}^{p},B_{i}^{p}))/\sum_{p=1}^{n}s(A_{i}^{p}),\vspace{-1mm}$
where $n$ is the number of parameters in $A_{i}$, $s(A_{i}^{p})$ denotes the number of values in the $p$-th parameter in $A_{i}$, and $cos(A_{i}^{p},B_{i}^{p})$ is the cosine similarity between the $p$-th parameters of $A_{i}$ and $B_{i}$. A higher value indicates a high similarity or equivalence between two Transformer layers.
Models with different embedding sizes. Checking equivalence among model components with different embedding sizes offers more flexibility in the sense that requests can potentially be routed to a smaller model component. It helps in balancing load and reducing latency, as processing through a smaller yet equivalent model component is faster.
Since size differences are common for models adapted from different foundation architectures, we focus on foundation model parameters. For example, the embedding sizes of 7B and 13B LLaMA are 4096 and 5120, respectively. Parametric cosine similarity is no longer applicable here anymore. Resorting to methods that enforce uniformity of dimensions, such as dimensional reduction, is impractical due to the extensive data requirements such processes entail.
(a) LLaMA 7B and LLaMA 13B.
(b) LLaMA 13B and LLaMA 33B.
Figure 10: Output cosine similarity Transformer layer groups of two models.
In BlockLLM, we leverage the output vocabulary probabilities as a means to ascertain such equivalence. By feeding the layers with the same data, we convert the output of each Transformer layer into vocabulary probabilities and compute the cosine similarity of the probability as an indicator of the degree of equivalence. In Figure 10, we have shown the output cosine similarity of Transformer Layer groups partitioned from different-sized models. The average similarity is 0.9841.

### v2

@@ Abstract
The increasing demand for Large Language Models (LLMs) across various applications has led to a significant shift in the design of deep learning serving systems. Deploying LLMs, particularly in multi-tenant environments, poses substantial challenges due to their high computational and memory demands. We introduce BlockLLM, a serving system that leverages component sharing among fine-tuned LLM models to provide an efficient and flexible solution for LLM workloads. BlockLLM partitions models into finer-grained blocks, enabling the reuse of model components and independent provisioning to improve computation efficiency. BlockLLM comprises an offline block zoo for storing blocks and an online system to serve requests through chains of blocks. It offers multi-fold flexibilities: (1) Adaptive assembly of blocks on-the-fly through equivalence evaluation among blocks in the zoo; (2) Per-block batch size configuration and best-effort KV cache coordination at the individual block level; (3) Speculative execution and locality-aware block placement to reduce communication costs from dynamic block resource allocation. Our evaluation shows that BlockLLM reduces memory and storage footprints and improves computational efficiency, outperforming existing serving approach in 95%ile latency and GPU utilization by 33.5% and 20.1%, respectively, with minimal impact on accuracy.
@@ 1 Introduction
The rise of Large Language Models (LLMs) marks a transformative milestone in deep learning. Their unprecedented abilities in natural language processing [36], from language translation [10] to question-answering [1] and complex reasoning [27, 49], are reshaping technological frontiers. Yet, the strength of LLMs—rooted in their vast parameter spaces—comes at the price of significant computation costs [9]. The deployment of LLMs is an arduous endeavor, often necessitating the use of hundreds or even thousands of extremely expensive computation devices like GPUs, thus imposing great challenges on unleashing LLM’s full potential [45, 6].
@@
Fine-tuning recently has emerged as a critical technique to efficiently adapt foundation models to handle various downstream tasks across different domains [21, 14]. Fine-tuning updates existing parameters or introduces additional parameters using domain-specific data. As LLMs become ubiquitous, the challenge of serving these fine-tuned models from a multi-tenant cluster—whether in a private cloud for first-party services or a public cloud for third-party users—has become increasingly significant.
@@
Recent serving systems and approaches focus on model-internal optimization, such as accelerating matrix computations [13, 2], minimizing device memory fragmentation [25], and enabling parallel decoding in the auto-regressive token generation process [16, 34]. While these advancements significantly accelerate inference of individual LLMs, the challenge remains to serve multiple tenants’ fine-tuned models in a shared cluster. Each LLM requires substantial resources, and the hardware cost makes it prohibitive to liberally expand the cluster to add dedicated resources for each tenant. Thus, the question is: how can LLM serving meet the latency goals of tenants’ fine-tuned models while also ensuring high throughput and optimal utilization cluster-wide?
@@
We introduce BlockLLM, an efficient and flexible serving system tailored for fine-tuned LLMs in a multi-tenant cluster. BlockLLM enables LLM model component reuse by breaking down an LLM into smaller and shareable components which we call “blocks”. The central observation forming the basis for BlockLLM is that fine-tuning offers the opportunity for sharing. Parameter-efficient fine-tuned models have an inherent modular architecture where parameters from specific layers are added or altered. Our observations with full-parameter fine-tuned models indicate that certain model components produce outputs with high similarity, suggesting their potential for sharing as well. Block sharing reduces both memory and storage demands. Furthermore, by reallocating memory—previously consumed by redundant parameters—towards supporting larger data batch size, BlockLLM achieves higher computation efficiency and overall throughput.
@@
BlockLLM consists of an offline “block zoo”—a repository of LLM blocks, and an online serving system capable of dynamically scheduling blocks to serve various applications (Figure 1). BlockLLM brings three immediate opportunities. (1) It can support adaptive serving, which enables on-the-fly assembly of blocks for each request, moving beyond the traditional static chain of blocks. (2) Each block can be individually configured, allowing for customized batch sizes and autonomous handling of queued requests. (3) BlockLLM enables dynamic resource allocation, with each block capable of independently scaling and being placed without the constraints of model-level boundaries.
@@
First, to ascertain the equivalence of blocks, we compare the output distribution of vocabulary probabilities, which serves as an indicator for functional equivalence between transformer layers. Furthermore, we investigate methods to facilitate the reuse and sharing of LLM components while minimizing any potential impact on accuracy. Inspired by the techniques of merging computer vision models [38], we design a generalizable stitching block as an intermediary to route requests among these blocks so that the chain of blocks to serve each request can be adaptively adjusted online.
Second, to manage the KV cache generated by different requests, BlockLLM employs a principled approach to request dispatching. The per-block configuration introduces complexity in handling the KV cache, particularly when requests are assigned to devices lacking the corresponding cache. This scenario necessitates either cache communication or recalculation. We analyze the trade-off between recalculation and I/O costs to determine priorities for different candidate block instances, including those specified in the chain and its equivalent blocks. We adopt a best-effort KV coordination policy, prioritizing devices that already possess the KV cache when dispatching requests.
Third, to address the inherent latency costs associated with inter-block communication, BlockLLM integrates speculative execution for identified bottleneck blocks, enabling the processing of potential request pathways in advance. We use surrogate models to predict the outputs of blocks, facilitating computation ahead of actual output. We also implement a locality-aware strategy for block placement, positioning interdependent blocks in close proximity, preferably within the same server. This approach minimizes the frequency of costly inter-server communications and leverages the high-speed intra-server connections to expedite data transfer.
@@
  • We leverage the LLM fine-tuning characteristic and demonstrate the benefits of finer-grained LLM serving. This approach reduces memory and storage demands while optimizing computation resource utilization.
  • We construct a block zoo that stores LLMs in finer granularity to enable block reuse and establishes the equivalence among these blocks to facilitate adaptive online serving.
@@
  • We implement and evaluate BlockLLM in a 12-A100 cluster. BlockLLM achieves comparable median latency with traditional per-model provisioning and reduces the 95%ile latency by 33.5%. The overall average throughput is elevated by 1.71x. GPU SM efficiency is improved by 20.1%.
@@ 2.1 Background
LLM serving. LLM-based applications have grown to become major workloads in GPU clusters, creating significant serving challenges. Large storage is required to store the model checkpoints and engines, and numerous powerful GPUs with ample memory are required to load and execute the models. [44, 39]. The auto-regressive nature of LLMs necessitates caching KV matrices to avoid redundant recalculations [40]. Each request’s KV cache grows as the generation process continues. Moreover, when models are inevitably deployed in a distributed manner, a high-capacity network is indispensable to transfer the intermediate activations. Such a burden is particularly pronounced in multi-tenant clusters, which cater to diverse LLM applications, each with its dedicated models and performance objectives.
@@
2.2 Key Idea: Finer-grained Serving
To enhance the efficiency of LLM serving systems, we advocate a finer-grained approach. Rather than treating each model as an indivisible unit, we propose BlockLLM, which partitions LLMs into smaller components provisioned independently. These components are envisioned as “blocks”. We could maintain only one copy of a block even if it is used by multiple models. The immediate benefit is the decreased demand for memory and storage.
@@ 2.2 Key Idea: Finer-grained Serving
(a) LLaMA 7B and LLaMA 13B.
(b) LLaMA 7B and Vicuna 7B.
Figure 3: Output similarity between Transformer layer groups of two models.
PE models are ideal for this approach due to their inherent modular architecture. As shown in Figure 2, two models could be divided into five blocks, with blue blocks being shared components. Moreover, our analysis of FF models reveals that different blocks share highly similar outputs.
Figure 3 presents the output vocabulary distribution similarity of each Transformer layer between LLaMA 7B and LLaMA13B, as well LLaMA 7B and Vicuna 7B(FF from LLaMA), with an average similarity of 0.9927 and 0.998, respectively. This observation underpins the feasibility of reusing blocks across different FF models, thereby enhancing resource efficiency by allowing device memory to hold more critical data.
@@
O3: Dynamic resource allocation. BlockLLM enhances resource allocation by scaling each block independently. Blocks that are frequently accessed or computationally intensive can dynamically be allocated with more resources. It also liberates block placement from the constraints associated with monolithic model architectures. However, it does introduce a challenge in the form of increased communication costs. Effectively mitigating these costs is crucial to achieving truly versatile resource allocation within BlockLLM (5.2, 5.3).
@@ 2.3 Comparison with Existing Work
Existing serving systems for serving LLM workloads utilize the same techniques as traditional DNN serving. Each model operates within a dedicated engine and is provisioned as a monolithic unit. An alternative approach is multi-tasking. A typical example is S-LoRA [43], which exploits memory paging to pack multiple LoRAs within one engine. Another line of optimizations focuses on the model internals—like accelerating Attention layer computations (FlashAttention [13]), managing KV cache (PagedAttention [25] in vLLM), and efficiently scheduling varying sequence lengths of requests (Orca [51]). These efforts are complementary to BlockLLM. BlockLLM tackles the inefficiency of serving LLM workloads in a multi-tenant cluster by leveraging the characteristics of LLM fine-tuning. Block-granularity provisioning improves the serving efficiency from two perspectives.
@@
Table 1: Percentage of shared parameters of different PE techniques.
@@ 3 BlockLLM Design
In this section, we provide an overview of the BlockLLM’s architecture and workflow, along with a discussion of its three key design challenges.
@@ 3.1 System Overview
Workflow. When a new request arrives, the scheduler assigns a chain of blocks based on the application it belongs to. The scheduler initially determines whether there is an available instance of the first block in the chain. If available, the request is forwarded to the target block instance. Otherwise, the scheduler either identifies a suitable alternative among the existing equivalent block instances or deploys a new block instance on an available device. Upon receipt of the request, the device’s agent directs it to the designated block instance for processing. Then, once the execution is completed, the agent is responsible for directing the request through the subsequent block in the chain. The process ends with the completion of the request signified by an EOS token, whereupon the final agent relays the output back to the scheduler, thus concluding the response cycle of the request.
@@ 3.2 Design Challenges
C3: Mitigate communication overhead. Versatile resource allocation of finer-grained blocks provides several benefits for online serving, as mentioned in 2.2. Yet BlockLLM incurs additional non-negligible communication overhead. Specifically, each time a request batch is directed to a different device, it necessitates intra-server or inter-server communication. This communication might act as a bottleneck for the batch, directly contributing to increased latency. The communication overhead also linearly grows with the number of block instances passed by a request batch during its lifecycle. Furthermore, the volume of data transferred escalates with the use of larger batch sizes per block. Figure 7 breaks down the latency components involved in generating one new token across various request batches. Without any optimization to the communication, for a request batch of size 32 and length 512, the computation constitutes 62.9% of the total latency, whereas the inter-server communication accounts for 36.4%.
@@ 4 Block Zoo
  • How to handle request routing between two equivalent blocks with different embedding sizes? (C1, 4.4)
@@ 4.1 Equivalence
PE models inherently possess equivalent components for reuse. We focus on verifying the equivalence among FF model components. Gvien that different models may produce outputs of varying sizes (e.g. the embedding sizes of the 7B and 13B LLaMA models are 4096 and 5120, respectively), we utilize output vocabulary probabilities to determine equivalence. By feeding the same data into the model, we transform the output of each Transformer layer into vocabulary probabilities and compute the cosine similarity of these probabilities as an indicator of equivalence. Figure 3 illustrates the output cosine similarity of layers partitioned from models of different sizes. Since LLaMA 7B and LLaMA 13B have different number of transfomer layers, we group 3 layers from LLaMA 7B and 4 layers from LLaMA 13B into a single block for comparison. The average similarity is 0.9841 for LLaMA 7B and LLaMA 13B, and 0.998 for LLaMA 7B and Vicuna 7B.
4.2 Accuracy Implications
(a) LLaMA 7B and Vicuna 7B.
(b) Vicuna 13B w/ 50% pruning.
Figure 10: Block Sharing Accuracy.
(a) LLaMA & Vicuna 7B switching.
(b) Vicuna 13B block switching.
Figure 11: Block Switching Accuracy.
Block Sharing. The output equivalence between foundational models and their FF models, as well as between differently-sized foundational models, suggests potential for dividing models into sharable blocks. However, naively sharing these blocks may impact inference quality, even with high output similarity.
Figure 10(a) shows how block sharing affects accuracy. We treated each decoding layer as a single block and tested various numbers of shared blocks at different layer positions. For instance, with four blocks, we used layers 0-3 of LLaMA 7B and layers 4-31 of Vicuna 7B, then layers 1-4 of LLaMA 7B and layers 0, 5-31 of Vicuna 7B, and so on. The figure shows significant accuracy impact when sharing occurs at initial layers, but minimal impact at middle and end layers.
This trend is consistent even with different block configurations. Figure 10(b) shows a similar pattern when applying 50% parameter pruning to blocks. Unlike the previous experiment, we only use blocks from Vicuna 13B, and applied 50% parameter pruning to selected blocks. The results indicate that pruning near the middle and end blocks has minimal accuracy impact, while initial blocks are less affected. Therefore, block sharing should be prioritized for middle or end components before initial ones. This phenomenon has also been observed in previous studies [33, 24, 11, 7], likely due to redundancy in later layers.
Block Switching. The design of BlockLLM allows a request to be processed by an arbitrary block at any point, which may result in a single request frequently switching between blocks and potentially affecting inference quality. Figure 11(a) shows the impact of frequent block switching on accuracy. We selected layer 16 as the starting point for block switching due to stable accuracy and treated $x$ subsequent layers as a single block. For each block after layer 16, we continuously switch blocks between two models. For instance, if $x=2$, a request will use layers 17-18 from Vicuna 7B, then switch to layers 19-20 of LLaMA 7B, then switch back to layers 21-22 of Vicuna 7B, and so on. We applied a similar switching strategy to 50% pruned Vicuna 7B blocks. Figures 11(a) and 11(b) show that frequent block switching near the end has minimal impact on accuracy. In practice, the switching does not occur so frequently due to exceessive communication overhead.